Wednesday, 30 June 2010


BT Wholesale are currently experiencing a major national fault meaning that users are unable to connect, or once connected may disconnect. They are working on this matter as a priority and will resolve it as quickly as possible.

We are expecting a BT Wholesale update at 10.30am.
We apologise to any customer affected by this.

Update from BT Wholesale:

This issue is still being worked on by BT engineers, the current fix ETA. is 13.00.

We will notify you of any further updates.

BT work continues, although most customer services have been restored.

We are awaiting a further update from BT at approx 15.00


BT Wholesale have confirmed this fault has now been resolved, and on their behalf we applogise to everyone who was affected by this.

Replacement hardware is on site and will be installed during an emergency maintenance window which is scheduled to take place between 00:00 and 06:00 tomorrow morning. Affected customers are expected to experience a disruption of service whilst this work is being carried out.


Apologies, we are seeing a reoccurence of the issue.
The maintenance work between 00:00 and 06:00 1st July should resolve this issue


Last nights work has resolved this issue.
Again we apologise for any inconvenience caused by this BT outage


Thursday, 10 June 2010

[CLOSED] - LLU Broadband Connectivity Issue

We are currently investigating issues with LLU connectivity.

This will only affect users on the LLU network.

We have just received the following report from our LLU provider:

"At approximately 10:30 (BST) today our LAC at THN stopped working and dropped all traffic routing through it. We are unable to contact it on its management interface and believe it may have crashed.

All traffic failed over to the LACs at TCH. The LACs at TCH are running at less than half load each, and are more than capable of handling all the network traffic on their own. However, we are now operating in a failover scenario so the network is at risk from a further failure.

We have despatched an engineer to THN to debug this locally and will update again later"

Once we receive the further update we will update this post.


Here is the update from our supplier:

"Our engineer on site was able to console into the LAC and establish that it had not crashed, but had lost arp layer communication.

We carried out various troubleshooting, including replacing SFPs and moving patches over to different router ports, but were unable to satisfactorily clear the fault or establish any cause from debugs. We have given the LAC a cold reboot and it is now performing normally.

We have brought the LAC back onto the network and allowed it to start taking some load again. We will be monitoring this device very closely now for any signs of an underlying software or hardware fault"

Please accept our apologies for any inconvenience caused.

Tuesday, 1 June 2010

[Closed] - Email Delays

Some customers are reporting delays in sending and receiving emails.

We are currently investigating this issue and will post our findings here.